OpenStudy (anonymous):

A=43degree and B=62dgree, and included side of c of length 14.6 a) find angle of C. =>75degree b)Find the length of side b.

OpenStudy (dumbcow):

use law of sines \[\frac{a}{\sin A} = \frac{b}{\sin B} = \frac{c}{\sin C}\] since you know c=14.6 , C=75, B=62 \[\frac{b}{\sin 62} = \frac{14.6}{\sin 75}\]